# Highly Fast Text Segmentation With Pairwise Markov Chains

### v2

@@ 1 Introduction
For over twenty years, HMCs have been strictly generalized to Pairwise Markov Chains (PMCs) [17], a family of models including HMCs. In PMCs, the “hidden" chain is not necessarily Markov, and the noise is modeled in a more correlated – and thus more informative – way. However, PMCs keep the same advantages as HMCs in regards to the hidden data estimation. In particular, the training and estimation tasks still have linear complexity. PMCs have especially been studied for image segmentation, with discrete hidden variables and continuous observations. It turns out that using PMCs instead of HMCs can divide the error rate by two [18] [19]. However, for specific reasons relative to language processing, which we will develop later, PMCs have never been applied for NLP tasks.
@@ 2.1 HMC - PMC
$\begin{split}p(z_{1:T})=&p(x_{1})p(y_{1}|x_{1})p(x_{2}|x_{1},y_{1})p(y_{2}|x_{2},x_{1},y_{1})\\ &...p(x_{T}|x_{T-1},y_{T-1})p(y_{T}|x_{T},x_{T-1},y_{T-1})\end{split}$ (3)
@@ 2.3 Parameter estimation
When $\text{card}(\Lambda_{X})$ or $\text{card}(\Omega_{Y})$ are huge, some patterns may not be observed in the training set, which implies that the corresponding estimation is zero. It is the case for NLP with $\Omega_{Y}$, the space of possible written words. To minimize this problem, especially in PMC, we accept to partially “downgrade" PMC to HMC when necessary. This original process is represented in figure 3. It consists of approximating the forward and backward probabilities of the PMC by the HMC ones when those of PMC equal 00.
@@ 3 Text segmentation tasks with Markov Chain models
  • $u(\omega)=1$ if the first letter of $\omega$ is up, 00 otherwise
  • $h(\omega)=1$ if $\omega$ has a hyphen, 00 otherwise
  • $f(\omega)=1$ if $\omega$ is the first word of the sentence, 00 otherwise
@@
It is important to note that the choice of the features is crucial as we cannot use arbitrary features with Markov chain models [23] [27] [28]. It preserves the model from a “second level” of unknown patterns, and it avoids having our new approximation of $b_{i}(k)$ to be equal to 00 too many times.
@@ 4 Experiments
CoNLL 2000 KW $1.94\%$ $1.72\%$ $\bm{1.27\%}$
[…]

### v3

@@ 1 Introduction
For over twenty years, HMCs have been strictly generalized to Pairwise Markov Chains (PMCs) [17] [18], a family of models including HMCs. In PMCs, the “hidden" chain is not necessarily Markov, and the noise is modeled in a more correlated – and thus more informative – way. However, PMCs keep the same advantages as HMCs in regards to the hidden data estimation. In particular, the training and estimation tasks still have linear complexity. PMCs have especially been studied for image segmentation, with discrete hidden variables and continuous observations. It turns out that using PMCs instead of HMCs can divide the error rate by two [19] [20]. However, for specific reasons relative to language processing, which we will develop later, PMCs have never been applied for NLP tasks.
@@ 2.1 HMC - PMC
$\begin{split}\hskip-28.45274ptp(z_{1:T})=&p(x_{1})p(y_{1}|x_{1})p(x_{2}|x_{1},y_{1})p(y_{2}|x_{2},x_{1},y_{1})\\ &...p(x_{T}|x_{T-1},y_{T-1})p(y_{T}|x_{T},x_{T-1},y_{T-1})\end{split}$ (3)
@@ 2.3 Parameter estimation
When $\text{card}(\Lambda_{X})$ or $\text{card}(\Omega_{Y})$ are huge, some patterns may not be observed in the training set, which implies that the corresponding estimation is zero. It is the case for NLP with $\Omega_{Y}$, the space of possible written words. To minimize this problem, especially in PMC, we accept to partially “downgrade" PMC to HMC when necessary. This original process is represented in figure 3. It consists of approximating the forward and backward probabilities of the PMC by the HMC ones when those of PMC equal 0.
@@ 3 Text segmentation tasks with Markov Chain models
  • $u(\omega)=1$ if the first letter of $\omega$ is up, 0 otherwise
  • $h(\omega)=1$ if $\omega$ has a hyphen, 0 otherwise
  • $f(\omega)=1$ if $\omega$ is the first word of the sentence, 0 otherwise
@@
It is important to note that the choice of the features is crucial as we cannot use arbitrary features with Markov chain models [24] [28] [29]. It preserves the model from a “second level” of unknown patterns, and it avoids having our new approximation of $b_{i}(k)$ to be equal to 0 too many times.
@@ 4 Experiments
CoNLL 2000 KW $1.94\%$ $\% 1.27\bm{1.27\%}$
CoNLL 2000 UW $16.54\%$ $\bm{12.47\%}$ $16.41\%$
CoNLL 2003 $5.29\%$ $\bm{4.28\%}$ $4.71\%$
CoNLL 2003 KW $4.03\%$ $\bm{3.21\%}$ $3.40\%$
CoNLL 2003 UW $15.30\%$ $\bm{12.79\%}$ $15.16\%$
UD English $8.13\%$ $\bm{6.95\%}$ $7.16\%$